# From EHRs to Patient Pathways: Scalable Modeling of Longitudinal Health Trajectories with LLMs

### v1

@@ 6 Conclusion
In this work, we introduced a novel approach to patient pathway modeling, leveraging a structured text-based representation of heterogeneous EHR data and a Masked Summarization Bottleneck to handle extended data categories and temporal contexts efficiently. By integrating diverse data modalities, including numerical, categorical, and free-text information, our model enables accurate next-step predictions and robust simulation of patient trajectories over extended time horizons. Our experiments demonstrate that this approach provides a scalable framework for holistic patient modeling. By evaluating both next-step and long-term simulations, we highlight the potential of our method to improve predictive healthcare and personalized care. This work establishes a foundation for further research into comprehensive, task-independent EHR modeling, paving the way for future advancements in personalized and predictive medicine.
@@ Acknowledgments
Appendix A Details on MIMIC-IV data
The MIMIC-IV [9] dataset can be obtained via PhysioNet [5, 8] after performing the necessary credential process and CITI Data or Specimens Only Research training under the PhysioNet Credentialed Health Data License 1.5.0. We use version 2.2. of the dataset provided at https://www.physionet.org/content/mimiciv/2.2/. In Table 6, we list all tables from the MIMIC-IV [9] dataset we include into our patient representation as well as the data categories included in the ICU chartevents table:
@@ Appendix A Details on MIMIC-IV data
Appendix B Qualitative example of Vital Sign Simulation
Figure 4 shows a qualitative example of the predicted and real development of different vital signs of a patient, predicted by the combined model EHR2Path-T+S.
@@ Appendix B Qualitative example of Vital Sign Simulation
Appendix C Example of our structured text representation
In Figure 5 and Figure 6, we show examples for the textual input and output representation for one time-point of a single patient.
@@ Appendix C Example of our structured text representation
Figure 6: Example of the expected output for the next hour.
@@
Appendix D Implementation Details
Our model is implemented in PyTorch and based on the Qwen-0.5b model [39], a compact LLM balancing capacity and efficiency, provided by unsloth [2] under the apache-2.0 license (https://huggingface.co/unsloth/Qwen2-0.5B-Instruct-bnb-4bit). We train using a next-token objective with LoRA [6] on structured text outputs (Section 3.1), optimized with AdamW ($1\times 10^{-4}$ learning rate). Training runs on a single NVIDIA A40 GPU (48 GB) with a batch size of 8 for one epoch, covering one million time points. The summary-based model trains in 1.5 days, while the 24h text-based model and the combined model require approximately 4.0 days. The summarization model, trained separately on one million section samples, also completes in 4.0 days. The text-based model has a 4000-token limit. In the summary-based model, each section of max. 5000 tokens is compressed into 8 summary embeddings, achieving up to 625x compression. For the summary+text model, we leverage the summary-based model as starting point and fine-tune it using curriculum learning to leverage additional text inputs. During fine-tuning we randomly drop the text or summary inputs in 1/3rd of the training samples each. This incentivizes the model to leverage both input types. At inference time our simulation takes on average 3.5 seconds per simulated hour of data on a single GPU. For the LOS Indicator ablation, we set the convergence limits as follows giving a 24 hour buffer for convergence: for ED Admission, we cap simulation at 65 steps (24 hours beyond the longest ED stay in our test set), while Discharge Diagnosis simulations, which always occurs within one time step, are constrained to 24 steps.
Appendix E Detailed Task Descriptions
@@
Discharge Diagnosis: Forecasts ICD categories at ED discharge, using all ED data up to that point. The LOS indicator is set to zero for direct prediction of disposition and ICD categories.
@@
Discharge Diagnosis: Forecasts ICD categories at hospital discharge, using all patient data up to that point. The LOS indicator is set to zero for direct prediction.
@@
Length-of-Stay (LOS): Predicts whether a patient will remain in the ICU for more than three days. Inputs include data up to a random time point, with simulation extending 72 hours or until discharge.

### v2

@@ 6 Conclusion
In this work, we introduced a novel approach to patient pathway modeling, proposing a structured text-based representation of heterogeneous EHR data and a Masked Summarization Bottleneck to handle extended data categories and temporal contexts efficiently. By integrating diverse data modalities, including numerical, categorical, and free-text information, our model enables accurate next-step predictions and robust simulation of patient trajectories over extended time horizons. Our experiments demonstrate that this approach provides a scalable framework for holistic patient modeling. By evaluating both next-step and long-term simulations, we highlight the potential of our method to improve predictive healthcare and personalized care. This work establishes a foundation for further research into comprehensive, task-independent EHR modeling, paving the way for future advancements in personalized and predictive medicine.
@@ Acknowledgments
Appendix A Appendix
A.1 Details on MIMIC-IV data
The MIMIC-IV Johnson et al. (2023b) dataset can be obtained via PhysioNet Goldberger et al. (2000); Johnson et al. (2023a) after performing the necessary credential process and CITI Data or Specimens Only Research training under the PhysioNet Credentialed Health Data License 1.5.0. We use version 2.2. of the dataset provided at https://www.physionet.org/content/mimiciv/2.2/, together with the corresponding versions of the MIMIC-IV-Note https://www.physionet.org/content/mimic-iv-note/2.2/ and the MIMIC-IV-ED dataset https://www.physionet.org/content/mimic-iv-ed/2.2/) (Johnson et al. (2021)). In table˜6, we list all tables from the MIMIC-IV (Johnson et al. (2023b)) dataset we include into our patient representation as well as the data categories included in the ICU chartevents table:
@@ A.1 Details on MIMIC-IV data
A.2 Qualitative example of Vital Sign Simulation
fig.˜6 shows a qualitative example of the predicted and real development of different vital signs of a patient, predicted by the combined model EHR2Path-S+T.
@@ A.2 Qualitative example of Vital Sign Simulation
A.3 Example of our structured text representation
In fig.˜7 and fig.˜8, we show examples for the textual input and output representation for one time-point of a single patient.
@@ A.3 Example of our structured text representation
(a) Example of an expected output for the next hour during the stay.
@@
(b) Example of an expected output for the next hour at the end of the stay.
Hospital Stay’:
Disposition’: DISCHARGED
ICD categories’: respiratory;pregnancy;congenital
Figure 8: Examples of an expected output for the next hour during the stay (top) and at the end of the stay (bottom).
A.4 Implementation Details
Our model is implemented in PyTorch and based on the Qwen-0.5b model Yang et al. (2024), a compact LLM balancing capacity and efficiency, provided by unsloth Daniel Han & team (2023) under the apache-2.0 license (https://huggingface.co/unsloth/Qwen2-0.5B-Instruct-bnb-4bit). We train using a next-token objective with LoRA Hu et al. (2022) on structured text outputs (section˜3.1), optimized with AdamW ($1\times 10^{-4}$ learning rate). At inference, we use the default Qwen decoding parameters ($temperature=0.7,top_{k}=20,top_{p}=0.8$). Training runs on a single NVIDIA A40 GPU (48 GB) with a batch size of 8 for one epoch, covering one million time points. The summary-based model trains in 1.5 days, while the 24h text-based model and the combined model require approximately 4.0 days. The summarization model, trained separately on one million section samples, also completes in 4.0 days. The text-based model has a 4000-token limit. In the summary-based model, each section of max. 5000 tokens is compressed into 8 summary embeddings, achieving up to 625x compression. For the summary+text model, we leverage the summary-based model as starting point and fine-tune it using curriculum learning to leverage additional text inputs. During fine-tuning we randomly drop the text or summary inputs in 1/3rd of the training samples each. This incentivizes the model to leverage both input types. At inference time our simulation takes on average 3.5 seconds per simulated hour of data on a single GPU. For all baselines, we use the original training and inference parameters. For all models, results are reported from a single run with bootstrapped 95% confidence intervals.
A.5 Detailed Task Descriptions
@@
Discharge Diagnosis: Forecasts ICD categories at ED discharge (multi-label prediction over 18 categories as defined by Slee (1978)), using all ED data up to that point. The LOS indicator is set to zero for direct prediction of disposition and ICD categories.
@@
Discharge Diagnosis: Forecasts ICD categories at hospital discharge (multi-label prediction over 18 categories as defined by Slee (1978)), using all patient data up to that point. The LOS indicator is set to zero for direct prediction.
@@
Imminent Discharge: Predicts whether a patient will be discharged from the ICU within 3 days. Inputs include data up to a random time point, with simulation extending 72 hours or until discharge.
Table 7 summarizes input window, simulation horizon, and label definition for each task.
Task Task Type Input
Window
Gap Window Output
Window
Rolling/Static
ED
Vital Signs
Development up to random ED timepoint 1h 24h or until ED stay end Rolling
ED
Admission
Outcome up to ED
admission
ED stay end Static
ED Discharge Diagnosis Outcome up to ED
discharge
Direct prediction Static
Hospital
Medications
Development up to random hospital
timepoint
1h 24h or until hospital discharge Rolling
Hospital
Lab Values
Development up to random hospital
timepoint
1h 24h or until hospital discharge Rolling
Hospital
Discharge
Diagnosis
Outcome up to discharge Direct prediction Static
ICU
Vital Signs
Development up to random ICU timepoint 1h 24h or until ICU discharge Rolling
ICU Inputs Development up to random ICU timepoint 1h 24h or until ICU discharge Rolling
ICU Imminent Mortality Outcome up to random ICU timepoint 1h within 24h Rolling
ICU Imminent Discharge Outcome up to random ICU timepoint 1h within 3 days Rolling
Table 7: Details of task definitions, including task type, input/ gap / output window and if the task is rolling or static.